OpenStudy (anonymous):

How do you solve ln(5x-1)=ln6-ln(x-6)

OpenStudy (anonymous):

you can use the property of log that: log(a) - log(b) = log(a/b) ln(5x - 1) = ln(6/(x-6)) take e of both sides 5x - 1 = 6/(x-6) (5x - 1)(x - 6) = 6. you can expand the left side and then find the roots with the quadratic formula. let me know if i you want me to solve for x

OpenStudy (anonymous):

other approach: log(a) + log(b) = log(ab). bring the ln(x - 6) to the left. e both sides

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

i'm kind of lost on what you do when you get to the step (5x-1)(x-6)=6

OpenStudy (anonymous):

you can expand by multiplying throughout. 5x(x) + 5x(-6) + (-1)x + (-1)(-6) = 6 5x^2 - 31x = 0

OpenStudy (anonymous):

maybe you've heard the acronym FOIL

OpenStudy (anonymous):

oh yeah i remember that part! so after that do i divide both sides by 5 or do something else?

OpenStudy (anonymous):

you can. I would factor out x. x(5x -31) = 0 this equation is true when x = 0 and when (5x - 31) = 0

OpenStudy (anonymous):

so is that the end result or do i have to do more? sorry to keep bothering you!

OpenStudy (anonymous):

it's not a bother. i'm here to help people :) i guess the questions wants you to solve for x. the answer would b x = 0 and x = 31/5 (obtained from 5x-31 = 0)

OpenStudy (anonymous):

thank you so much! :)